given f(x)=6x+1, g(x)=8x+7 find f(g(x))
Answer by jim_thompson5910(35256)

given f(x)=6x+1, g(x)=8x+7 find f(g(x))
-------------------------------------------------------

f(x)=6x+1

f(x)=6( x )+1

f(g(x))=6( g(x) )+1

f(g(x))=6( 8x+7 )+1

f(g(x))=48x+42+1

f(g(x))=48x+43


Final Answer: f(g(x))=48x+43